# The Mind in the Machine: A Survey of Incorporating Psychological Theories in LLMs

### v1

@@ 6.3 What Debates Loom at the NLP–Psychology Intersection, and Where Next?
A recurring question is whether human psychology can be directly mapped to LLMs without distortion (Löhn et al., 2024). Below, we highlight key controversies at this boundary; see Appendix D for an extended discussion. These challenges motivate new recommendations and highlight open directions for cross-disciplinary exploration.
@@
A core tension is the mismatch between psychological terminology and their NLP usage. For example, attention in psychology means selective mental focus, but in transformers it is a token weighting mechanism without cognitive awareness (Lindsay, 2020), leading to misleading attributions of intentionality. Similarly, memory in psychology entails structured encoding and recall, whereas in LLMs it typically refers to context windows or parameters. Such anthropomorphic language is increasingly prevalent and shapes public and scholarly assumptions about LLMs, as recent studies show rising human-like descriptors (Ibrahim & Cheng, 2025). This calls for disentangling metaphor from mechanism through a precise cross-disciplinary lexicon, preventing both oversimplification and over-anthropomorphization – an underexplored but crucial research challenge.
Theoretical Discrepancies in Use of Psychology
Beyond terminology, deeper theoretical mismatches arise when the NLP community adopts outdated or disputed concepts from psychology. For instance, predictive coding (Rao & Ballard, 1999) is used to analogize LLMs’ next-token prediction, although current research emphasizes hierarchical, multi-scale brain mechanisms (Antonello & Huth, 2024; Caucheteux et al., 2023). Likewise, folk-psychological typologies like MBTI persist in LLM applications despite its criticized validity and reliability (Pittenger, 1993; McCrae & Costa Jr, 1989).
Working memory (Baddeley & Hitch, 1974a) illustrates another gap: LLM ’memory’ modules (Kang et al., 2024; Li et al., 2023) do not replicate human constraints, prompting questions about whether AI should emulate human cognitive limits or exceed them for performance gains. Behavioral psychology faces similar critiques (Miller, 2003; Flavell et al., 2022), as RLHF often focuses on reward optimization (Ouyang et al., 2022; Rafailov et al., 2023; Ramesh et al., 2024), neglecting internal states and risking reward hacking (Skalse et al., 2022; Krakovna, 2020). Broader debates remain over whether LLMs truly “understand” language or function as “stochastic parrots” (Ambridge & Blything, 2024; Park et al., 2024).
@@
A parallel controversy involves personality: some studies find stable simulated traits (Sorokovikova et al., 2024; Huang et al., 2024), while others reveal variability under different prompt conditions (Gupta et al., 2024; Shu et al., 2024), raising questions about inherent vs. mimicked personas (Tseng et al., 2024). These debates underscore the need for a systematic, theory-driven framework beyond surface metrics, guiding more faithful replication of human cognition and behavior in LLMs.
@@ 7 Conclusions
We systematically review how psychology can ground LLM innovation in both past and future across key subfields: cognitive, developmental, behavioral, social, personality, and psycholinguistic. We examine how psychological theories inform each stage of LLM development, revealing both meaningful connections across domains and critical points of tension, which we explore through discussion to help bridge interdisciplinary gaps. We hope this survey sparks reflection, and inspires future work to continue integrating psychological perspectives into NLP in meaningful and impactful ways.
@@
Our survey primarily focuses on literature within NLP, particularly in how personality is modeled, evaluated, and leveraged in LLMs. As a result, we do not extensively cover research from psychology and cognitive sciences that might offer deeper theoretical insights into human-like behaviors in AI. This limitation may exclude valuable methodologies or perspectives that could enhance personality evaluation frameworks for LLMs. We encourage future surveys to integrate findings from psychology and linguistics to bridge theoretical foundations with computational approaches, fostering a more comprehensive understanding of personality in AI systems.
While our survey advocates for a deeper integration of psychology into LLM design, we also caution against the ethical risks posed by overuse or misapplication of psychological principles. A concrete example is operant conditioning (Skinner, 1957), which describes how behavior can be shaped by consequences. Applied to LLMs – for instance, through timely, gratifying feedback to reinforce engagement – these mechanisms can be beneficial in contexts like language learning or motivation. However, reinforcement schedules such as variable ratio or interval rewards may unintentionally condition users to engage compulsively, raising the risk of manipulative design. This presents a key ethical limitation: distinguishing between genuinely supportive interactions and those that encourage excessive use is inherently difficult. To address this, we emphasize the need for transparent disclosure of reinforcement mechanisms and the establishment of clear ethical guidelines by professional communities. These safeguards are essential to ensure that psychological insights enhance user well-being without enabling exploitative practices.
Appendix
@@ Appendix A Psychology Theories
{NiceTabular}
C65pt—c—C80pt—C250pt—c Psych Area Sub Area Theory Definition Explored
[…]
Scaffolding Theory Learning is enhanced through gradually challenging interactions with appropriate guidance (Park & Reuter-Lorenz, 2009)
Incremental Numerical Understanding Numerical concepts are gradually acquired through structured exposure and experience (Piaget, 2013)
Zone of Proximal Development Optimal learning occurs in the gap between what a learner can do independently and with assistance (Wertsch, 1988)
[…]
Ecological Validity Emphasizes real-world data and environments to mimic natural cognitive development (Schmuckler, 2001)
[…]
Top-down and Bottom-up Processing Distinguishes between concept-driven (top-down) and data-driven (bottom-up) perceptual processing (Gregory, 1997)
Predictive Coding Anticipatory processing based on prior knowledge and prediction of expected inputs (Rao & Ballard, 1999)
[…]
Long-term Memory System for storing information over extended periods through semantic organization (Tulving et al., 1972)
Hippocampal Indexing Theory Views the hippocampus as a pointer to neocortical memory representations (Teyler & DiScenna, 1986)
[…]
Theory of Mind The ability to attribute mental states to oneself and others and understand others may have different beliefs (Baron-Cohen et al., 1985)
Schema Theory Knowledge is organized into interconnected patterns that guide processing and interpretation of new information (Anderson & Pearson, 1984)
[…]
Operant Conditioning Behavior is strengthened or weakened by consequences such as rewards or punishments (Skinner, 1957; 1938)
Thorndike’s Law of Effect Behaviors followed by satisfying outcomes are more likely to be repeated in the future (Thorndike, 1927)
Premack Principle A preferred activity can reinforce a less preferred one if access is contingent (Premack, 1959)
Table 1: Representative developmental, cognitive, and behavioral psychology theories by sub-area. In the “Explored” column,
indicates multiple surveyed works,
indicates fewer than three, and
indicates that none emerged in our survey (i.e., not yet substantially explored).
{NiceTabular}
C65pt—c—C80pt—C250pt—c Psych Area Sub Area Theory Definition Explored
[…]
Dual-Process Theory Differentiates between fast, intuitive (System 1) and slow, deliberate (System 2) reasoning (Kahneman, 2011)
Theory of Mind (ToM) How individuals understand and attribute mental states to others (Baron-Cohen et al., 1985)
[…]
4-1Social Influence Social Impact Theory The magnitude of social influence depends on the strength, immediacy, and number of sources (Latané, 1981)
Conformity Theories Explore how group pressure can alter individual judgments (Asch, 2016)
Obedience Theories Demonstrate how authority influences behavior, highlighting conditions under which individuals comply (Milgram, 1963)
Persuasion Models Explain how messages processed via central or peripheral routes can lead to attitude change (Petty & Cacioppo, 2012)
[…]
2-1Group Dynamics Groupthink Examines how the desire for conformity and group cohesion can lead to flawed decision-making and suppression of dissenting opinions (Janis, 1972)
Social Facilitation and Social Loafing Investigates how the presence of others can enhance performance on simple tasks or reduce effort in collective work (Zajonc, 1965; Latané et al., 1979)
[…]
4-1Attitude Change Cognitive Dissonance Theory Explains how inconsistencies between beliefs or behaviors create discomfort, prompting attitude change to restore consistency (Morvan & O’Connor, 2017)
Elaboration Likelihood Model (ELM) Proposes that persuasion occurs via a central route (deliberate processing) or a peripheral route (heuristic processing), depending on the recipient’s motivation and capacity (Petty & Cacioppo, 2012)
Balance Theory Suggests that individuals strive for consistency among their attitudes and relationships, adjusting beliefs to maintain cognitive harmony (Heider, 1946)
Inoculation Theory Posits that exposure to weak counterarguments can strengthen resistance to persuasion by preemptively activating defensive mechanisms (McGuire, 1964)
[…]
5-1Self and Identity Self-Reflection Defines the process of introspection, with attention placed on the self-concept (Phillips, 2020)
Self-Perception Theory Explains how individuals infer their internal states by observing their own behavior (Bem, 1972)
Social Identity Theory Posits that group membership shapes self-concept and influences intergroup behavior (Tajfel, 1979)
Self-Categorization Theory Expands on social identity theory, describing how individuals classify themselves and others into social groups, shaping social norms (Maines, 1989)
Self-Affirmation Theory Suggests that individuals are motivated to maintain their self-integrity when faced with threats to their self-concept (Steele, 1988)
Table 2: Representative social psychology theories by sub-area. In the “Explored” column,
indicates multiple surveyed works,
indicates fewer than three, and
indicates that none emerged in our survey (i.e., not yet substantially explored).
{NiceTabular}
C65pt—c—C80pt—C250pt—c Psych Area Sub Area Theory Definition Explored
[…]
Myers-Briggs Type Indicator (MBTI) Classifies individuals into 16 personality types based on four dichotomies (e.g., Introversion vs. Extraversion) (Myers & Myers, 1995). While widely used, MBTI has been criticized for lacking empirical validity, reliability, and independence between its categories. (Pittenger, 1993)
Eysenck Personality Questionnaire-Revised (EPQR-A) Contains a 24-item personality test that measures extraversion, neuroticism, psychoticism, and social desirability. (Eysenck & Eysenck, 1984)
[…]
5-1Personality Theories Humanistic Theory Emphasizes free will, personal growth, and self-actualization. This perspective focuses on individuals’ subjective experiences and their drive to achieve their full potential. (Stefaroi, 2015)
Psychoanalytic Theory Originating from Freud, this theory conceptualizes personality as the dynamic interplay between the id, ego, and superego, with unconscious processes playing a central role in shaping behavior. (Scharff et al., 2013)
Behaviorist Theory Views personality as a set of learned responses shaped by environmental reinforcements and punishments. This perspective, pioneered by figures like Skinner and Watson, rejects internal mental states in favor of observable behaviors. (Pierce & Cheney, 2008)
Social Cognitive Theory Highlights the role of cognitive processes in personality, emphasizing how expectations, beliefs, and observational learning shape behavior. (Spielman et al., 2024)
Trait Theory Focuses on identifying and measuring stable personality traits that influence behavior across different contexts. (Cartwright, 1979)
[…]
Usage-Based Theory Emphasizes the role of social interaction and cognitive processes in language learning, rather than innate universal grammatical structures (Tomasello, 2005)
[…]
Constraint-Based Models Language processing is an interactive, probabilistic process where multiple sources of information simultaneously contribute to understanding, rather than following a strict, sequential parsing approach (MacDonald et al., 1994)
Good-Enough Processing Proposes that humans comprehend language through approximate, semantically-focused representations that capture the core meaning rather than constructing syntactically perfect linguistic interpretations (Ferreira & Patson, 2007)
Construction-Integration Model Describes text comprehension as a two-stage process where readers first generate multiple, loosely connected propositions and then systematically filter and integrate them into a coherent, meaningful understanding. (Kintsch, 1988)
[…]
Interactive Two-Step Model An interactive, probabilistic process of lexical selection and phonological encoding, where multiple linguistic levels simultaneously influence each other during speech generation (Goldrick & Rapp, 2007)
Table 3: Representative personality psychology and psycholinguistics theories by sub-area. In the “Explored” column,
indicates multiple surveyed works,
indicates fewer than three, and
indicates that none emerged in our survey (i.e., not yet substantially explored).
In Tables A, A, and A, we summarize representative theories in each psychological domain and note whether they have been leveraged in LLM development.
Appendix B Extended Discussion on Reinforement Learning from Human Feedback (RLHF)
B.1 Operant Conditioning in RLHF
During RLHF fine-tuning, the model (agent) generates responses while a learned reward function $R(x)$, often a neural network trained on preference data, assigns scores to candidate outputs $x$. These scores proxy for human judgment and guide policy updates to reinforce higher-reward behaviors. For instance, Ouyang et al. (2022) trains InstructGPT via Proximal Policy Optimization (Schulman et al., 2017): responses deemed more helpful or accurate by human evaluators receive greater reward, whereas undesirable or incorrect outputs face penalization. Unlike purely exploration-based RL methods, this arrangement leverages human insight to provide a more precise learning signal; however, success relies on careful and consistent reward modeling that captures subtle human values.
B.2 Modeling Human Preferences as a Reward Function
Although extensive work has been conducted in RLHF, here we primarily highlight recent approaches or methodologies explicitly grounded in psychological theories. Building robust reward functions from heterogeneous or ambiguous feedback remains a core challenge in RLHF. Early foundational frameworks (Christiano et al., 2017; Stiennon et al., 2022) laid essential groundwork for converting human judgments into usable reward signals, drawing implicitly from principles of Operant Conditioning Theory. More recent advancements explicitly target improvements in stability, scalability, and fairness, addressing issues arising from the inherent variability and complexity of human preferences.
Rafailov et al. (2023) introduced Direct Preference Optimization (DPO), simplifying preference integration by directly optimizing the policy through a closed-form solution, thus removing the need for explicit intermediate reward modeling. Extending these efforts toward equitable alignment, Ramesh et al. (2024) proposed Group Robust Preference Optimization (GRPO), ensuring robustly aligned outcomes across diverse demographic groups, addressing biases commonly observed in human-driven reward processes.
Further refinements emphasize enhancing alignment accuracy through psychological considerations. For instance, Contrastive Preference Learning (Hejna et al., 2024) utilizes regret-based losses inspired by behavioral economics, facilitating stable off-policy learning without conventional RL techniques. Distributional Preference Learning (Siththaranjan et al., 2024) aligns reward modeling more closely with human cognitive patterns by capturing human values as probability distributions rather than point estimates. Variational Preference Learning (VPL) (Poddar et al., 2024) further integrates psychological realism, introducing latent-variable modeling to personalize RLHF, reflecting variability in individual user preferences rather than imposing a universal reward structure.
These advancements collectively illustrate how psychological theory, particularly Operant Conditioning Theory, continues to shape and inspire sophisticated techniques for reliably aligning LLM behavior with nuanced human values.
B.3 Reinforcement Schedules and Feedback Frequency
In early RLHF, feedback is typically sparse — a single scalar reward per output — which causes a credit assignment problem: the model can’t tell which parts of the output led to the reward. This is similar to delayed feedback in animal learning, which slows progress. Psychology shows that immediate and frequent reinforcement improves learning. Similarly, recent RLHF methods provide dense, token-level feedback (e.g., from a critic model), which improves sample efficiency and training stability. To address this, Cao et al. (2024) propose LLM self-critique, a method that uses a secondary model to provide dense, token-level feedback during generation. This simulates a continuous reinforcement schedule, analogous to real-time feedback in behavioral training, and leads to more stable and efficient learning. Another factor is how often feedback is given: continuous vs. partial reinforcement. While human feedback is often sparse due to cost, using AI feedback models (like RLAIF, will discuss later) allows for more frequent feedback. Even with limited human scores, techniques like credit assignment can distribute reward across the output.
B.4 Reward Prediction Errors as a Learning Driver
At the heart of reinforcement learning lies the concept of reward prediction error (RPE), which arises when there is a discrepancy between an agent’s expected reward and the reward it actually receives, prompting adjustments and driving learning (Sutton & Barto, 2018). This mechanism closely parallels dopaminergic signaling in animal brains, where dopamine neurons respond strongly to unexpected rewards or punishments, effectively reinforcing behaviors associated with positive surprises or reducing those linked to disappointments (Schultz, 1998). In RLHF, reward prediction errors similarly guide model updates; each model output receives a score from a reward model trained on human preferences, and deviations between these scores and the model’s predicted rewards are used to adjust behavior. However, simplistic or flawed reward models can lead to ”reward hacking,” where the model exploits blind spots in the reward function rather than genuinely aligning with human values (Amodei et al., 2016). Introducing variability in reward signals can encourage exploration and mitigate premature convergence on suboptimal strategies (Dayan & Daw, 2008). To address reward hacking and reward-model inconsistencies, recent approaches have formulated RLHF as a constrained Markov decision process with dynamic weighting (Moskovitz et al., 2024), introduced information-theoretic regularization techniques (InfoRM) (Miao et al., 2024b), and proposed methods such as ConvexDA and reward fusion to stabilize and enhance reward-model consistency (Shen et al., 2024).
B.5 Implications for Bias, Alignment, and Reward Modeling
Employing these behavioral principles may improve how well RLHF handles biases and achieves robust alignment. For instance, diverse trainers and variable scenarios can prevent conditioning bias, where the model overfits to a narrow segment of human preferences (Sheng et al., 2019). Moreover, shaping and multi-dimensional reward functions can address multiple alignment goals simultaneously (e.g., factual accuracy and polite style), limiting reward hacking.
At the same time, grounding RLHF in behavioral theory highlights persistent pitfalls. Models still lack an intrinsic understanding of human values, and an imprecise reward signal can reinforce superficial behaviors. To mitigate these risks, a cycle of model auditing, reward model refinement, and re-training can mirror how animal trainers continually adjust reinforcement to avoid unwanted side effects.
Appendix C Persona-Inspired Dialogue Generation
Personality has also inspired improvements truthfulness, response grounding, and broader alignments. Zhang et al. (2024d) introduced Self-Contrast to enhance internal consistency, and Joshi et al. (2024) proposed the Persona Hypothesis, linking truthfulness to pretraining structure. Kim et al. (2024) introduced PANDA to reduce persona overuse in dialogue. Zhang et al. (2024d) introduced a reflection-based technique to reduce internal inconsistencies. Joshi et al. (2024) proposed the Persona Hypothesis, arguing that LLMs encode truthful and untruthful personas from their training distribution. Kim et al. (2024) addressed the overuse of persona cues to improve contextual appropriateness. Persona-guided generation has been applied to emotionally supportive role-play settings (Ye et al., 2025; Chen et al., 2025b).
Appendix D Extended Discussion on Debates over NLP-Psychology Intersection
A recurring theme is whether human psychology can be naively mapped onto LLM behavior without distortion (Löhn et al., 2024). Therefore, in this section, we discuss several major points of contention at this interdisciplinary boundary. These issues motivate a set of recommendations and highlight open directions for future cross-disciplinary research.
Terminology Mismatches
One key issue is the mismatch in terminology and the anthropomorphization of technical concepts. Terms like attention, memory, and “understanding” have specific meanings in psychology that differ from their usage in NLP. For instance, attention in psychology refers to selective mental focus and executive control, whereas in transformers models, it is a mathematical mechanism for weighting tokens – without cognitive awareness (Lindsay, 2020). This divergence can lead to misleading interpretations, such as assuming models exhibit intentional focus when they merely perform matrix operations. Similar misalignments exist for terms like memory (which in psychology implies a structured encoding and recall process, versus an LLM’s context window or weight parameters) and expressions such as “knows” or “thinks.”
Such anthropomorphic language is increasingly prevalent and shapes public and scholarly assumptions about LLMs. Recent analyses have found a growing prevalence of human-like descriptors for LLM behavior, raising calls to carefully disentangle metaphor from mechanism (Ibrahim & Cheng, 2025). An open research direction is developing a more precise cross-disciplinary lexicon: how can we describe model behaviors in ways that neither oversimplify the psychology nor over-anthropomorphize the engineering? Improving interdisciplinary communication by explicitly defining terms and drawing careful analogies remains an important but under-addressed challenge.
Theoretical Discrepancies in Use of Psychology
Beyond terminology, discrepancies arise in the adoption of psychological theories within NLP research. Sometimes, NLP integrates concepts from psychology that are outdated or contested in their original fields. For instance, predictive coding, which proposes that the brain continuously anticipates sensory input and updates via prediction errors (Rao & Ballard, 1999), is often used as a metaphor for LLMs’ next-token prediction. However, contemporary studies emphasize that brain prediction operates across hierarchical and multi-scale structures (Antonello & Huth, 2024; Caucheteux et al., 2023), cautioning against simplistic analogies that risk misrepresenting the theory.
Another example is the lingering use of folk-psychological typologies like the MBTI in some LLM studies. Despite its cultural popularity, MBTI has faced substantial criticism for poor validity and reliability (Pittenger, 1993). It classifies personality into 16 types based on Jungian dichotomies; however, research indicates these categories lack stability and predictive power regarding behavior (McCrae & Costa Jr, 1989). Nonetheless, the ease of obtaining of MBTI-labeled data has led some NLP studies to treat these categories as definitive, highlighting a theoretical lag where NLP adopts psychological models that mainstream psychology has largely moved beyond.
Working memory presents another gap. While cognitive psychology and neuroscience characterize it by limited capacity and active attention control (Baddeley & Hitch, 1974a), LLM approximations – such as short-term retention modules (Kang et al., 2024) or memory mechanisms for external context (Li et al., 2023) – do not replicate these constraints. This raises questions: Should AI systems emulate human cognitive limitations to achieve more human-like reasoning, or should they leverage their capacity to surpass such constraints? If certain human limitations, like bounded memory, lead to desirable properties such as better interpretability or reduced distractions, might it be useful to impose similar limits on AI? These questions remain largely open.
Finally, a related debate concerns behavioral psychology. The field has been critiqued for ignoring cognitive processes (Miller, 2003) and internal mental states (Flavell et al., 2022) that drive the observed behaviors, limiting its explanatory power. With the critiques remaining, the superficial application of behavioral psychology is also evident in LLM research. For instance, RLHF draws from operant conditioning but largely focuses on optimizing rewards (Ouyang et al., 2022; Rafailov et al., 2023; Ramesh et al., 2024), often neglecting internal model states. Consequently, a flip-side of such optimization is reward hacking (Skalse et al., 2022), where models exploit shortcuts without meeting true objectives – mirroring human behavior under evaluative pressure (Krakovna, 2020). Deeper integration of cognitive psychology is needed to address these limitations in LLM design.
The debate over whether LLMs possess a true understanding of language or merely function as ”stochastic parrots” (Bender et al., 2021) remains ongoing. Linguists have largely been skeptical (Ambridge & Blything, 2024), arguing that language ability is inherently abstract and complex, extending beyond mere statistical pattern recognition. Park et al. (2024) connection between mathematical reasoning and high-level linguistic comprehension.
Evaluation and Validity Debates
Anoter central debate concerns how we evaluate LLMs on purportedly “psychological” abilities – and whether current tests measure what we assume. For example, advanced LLMs like GPT-4 perform well on traditional ToM tasks, solving around 75% of false-belief scenarios, comparable to a 6-year-old child (Kosinski, 2024; Strachan et al., 2024). Some interpret this as emergent ToM-like reasoning (Kosinski, 2024), but others caution that high performance may reflect surface-level pattern matching rather than genuine mental-state attribution. Researchers emphasize that correct answers do not imply mentalizing ability (Strachan et al., 2024), and minor prompt changes can significantly impair model performance (Shapira et al., 2024). This underscores the need for more rigorous, theory-grounded evaluations and clearer cross-disciplinary definitions.
A similar controversy surrounds personality modeling. Some studies suggest LLMs exhibit stable simulated personality traits (Sorokovikova et al., 2024; Huang et al., 2024), enabling consistent persona simulation across prompts. However, others show that LLM responses vary with prompt framing and response order, undermining test reliability (Gupta et al., 2024; Shu et al., 2024). Tseng et al. (2024) distinguish between role-playing (adopting assigned traits) and personalization (adapting to users), raising a fundamental question: do LLMs have inherent personalities, or merely mimic behavior? While LLMs can simulate personality, inconsistent assessments cast doubt on whether such traits are emergent or engineered – an open direction for future work.
In summary, these debates highlight the need for a systematic, theory-driven framework that goes beyond superficial performance metrics, thereby enhancing model interpretability and guiding the development of LLMs to more faithfully replicate the complexities of human cognition and behavior.

### v2

@@ 6.3 What Debates Loom at the NLP– Psychology Intersection, and Where Next?
A recurring question is whether human psychology can be directly mapped to LLMs without distortion (Löhn et al., 2024). Below, we highlight key controversies at this boundary; see Appendix LABEL:appendix:debates for an extended discussion. These challenges motivate new recommendations and highlight open directions for cross-disciplinary exploration.
@@
A core tension is the mismatch between psychological terminology and their NLP usage. For example, attention in psychology refers to selective mental focus, but in transformers it is a token weighting mechanism without cognitive awareness (Lindsay, 2020), leading to misleading attributions of intentionality. Similarly, memory in psychology entails structured encoding and recall, whereas in LLMs it typically refers to context windows or parameters.
Such anthropomorphic language is increasingly prevalent and shapes public and scholarly assumptions about LLMs, as recent studies show rising human-like descriptors (Ibrahim and Cheng, 2025). This calls for disentangling metaphor from mechanism through a precise cross-disciplinary lexicon, preventing both oversimplification and over-anthropomorphization – an underexplored but crucial research challenge.
Theoretical Discrepancies
Beyond terminology, deeper theoretical mismatches arise when the NLP community adopts outdated or disputed concepts from psychology. For instance, predictive coding (Rao and Ballard, 1999) is used to analogize LLMs’ next-token prediction, although current research emphasizes hierarchical, multi-scale brain mechanisms (Antonello and Huth, 2024; Caucheteux et al., 2023). Likewise, folk-psychological typologies like MBTI persist in LLM applications despite its criticized validity and reliability (McCrae and Costa Jr, 1989; Pittenger, 1993). Wagner et al. (2025) positions that ToM involves first deciding depth of mentalizing and then applying reasoning accordingly, yet most works focus only on the latter. Working memory (Baddeley and Hitch, 1974) illustrates another gap: LLM "memory" modules (Li et al., 2023; Kang et al., 2024) do not replicate human constraints, prompting questions about whether AI should emulate human cognitive limits or exceed them for performance gains. Behavioral psychology faces similar critiques (Flavell et al., 2022; Miller, 2003), as RLHF often focuses on reward optimization (Ouyang et al., 2022; Rafailov et al., 2023; Ramesh et al., 2024), neglecting internal states and risking reward hacking (Krakovna, 2020; Skalse et al., 2022). Broader debates remain over whether LLMs truly “understand” language or function as “stochastic parrots” (Ambridge and Blything, 2024; Park et al., 2024).
@@
A parallel controversy involves personality. Some studies find stable simulated personality traits (Sorokovikova et al., 2024; Huang et al., 2024), while others reveal variability under different prompt conditions (Gupta et al., 2024; Shu et al., 2024), raising questions about inherent vs. mimicked personas Tseng et al. (2024). These debates underscore the need for a systematic, theory-driven framework beyond surface metrics, guiding more faithful replication of human cognition and behavior in LLMs.
These debates also echo concerns raised within Psychologists’ Perspectives on LLMs, where scholars are critically examining the promises and limitations of applying human psychological constructs to LLMs. Psychologists are engaging LLMs both as tools for research and as objects of study, but they caution against straightforwardly mapping human psychological constructs onto models trained via next-token prediction on disembodied text. Method papers flag opportunities and pitfalls for using LLMs as participants, judges, or simulators, urging stronger validity checks and domain knowledge when importing instruments or constructs Abdurahman et al. (2024b). At the theoretical level, scholars argue we should be precise about what LLMs are meant to model and avoid inferring humanlike learning or cognition from surface similarities Blank (2023); Cuskley et al. (2024). Empirically, comparative studies probe where model behavior diverges from humans on socially loaded tasks: LLMs can pass certain ToM-style probes while still differing on harder variants Strachan et al. (2024); Jones et al. (2024), and they sometimes amplify cognitive biases such as omission bias in moral judgment Cheung et al. (2025). Together, these works show an active psychology literature investigating LLMs’ promise and limits, and underscores why direct transfer of human psychological theory to LLMs requires careful construct validity, task design, and interpretation.
@@ 7 Conclusions
We systematically review how psychology can ground LLM innovation in both past and future across six key subfields. We examine how psychological theories inform each stage of LLM development, revealing both meaningful connections across domains and critical points of tension, which we explore through discussion to help bridge interdisciplinary gaps. We hope this survey sparks reflection, and inspires future work to continue integrating psychological perspectives into NLP in meaningful and impactful ways.
@@
Our review primarily focuses on literature within NLP, particularly in how personality is modeled, evaluated, and leveraged in LLMs. As a result, we do not extensively cover research from psychology and cognitive sciences that might offer deeper theoretical insights into human-like behaviors in AI. This limitation may exclude valuable methodologies or perspectives that could enhance personality evaluation frameworks for LLMs. We encourage future surveys to integrate findings from psychology and linguistics to bridge theoretical foundations with computational approaches, fostering a more comprehensive understanding of personality in AI systems. Some perspectives (e.g., embodied cognition, constructivism) cut across our categories and appear sporadically in current LLM work; we note these as cross-cutting threads and encourage future surveys to map them more fully across stages."
While our survey advocates for a deeper integration of psychology into LLM design, we also caution against the ethical risks posed by overuse or misapplication of psychological principles. A concrete example is operant conditioning (Skinner, 1957), which describes how behavior can be shaped by consequences. Applied to LLMs, these mechanisms can be beneficial in contexts like language learning or motivation. For instance, through timely, gratifying feedback to reinforce engagement. However, reinforcement schedules such as variable ratio or interval rewards may unintentionally condition users to engage compulsively, raising the risk of manipulative design. This presents a key ethical limitation: distinguishing between genuinely supportive interactions and those that encourage excessive use is inherently difficult. To address this, we emphasize the need for transparent disclosure of reinforcement mechanisms and the establishment of clear ethical guidelines by professional communities. These safeguards are essential to ensure that psychological insights enhance user well-being without enabling exploitative practices.
We also recognize longstanding critiques of psychology, for example, its emphasis on individual-level constructs, WEIRD sampling (overreliance on participants from Western, Educated, Industrialized, Rich, and Democratic societies in behavioral and social science research), and under-representation of social/structural forces. Our survey foregrounds psychology because many current LLM papers explicitly reference it; however, insights from linguistics and sociology are equally important for understanding language as a fundamentally social phenomenon. Future surveys that braid these traditions more tightly would provide a more complete view.
An important adjacent question is what LLMs do psychologically to users (e.g., persuasion, behavior change, support). We flag this as a crucial direction for future surveys that integrate psychology with HCI and communication research to evaluate user-level outcomes alongside model-level capabilities.
Acknowledgments
This work is in part supported by the funds provided by the National Science Foundation and by DoD OUSD (R&E) under Cooperative Agreement PHY-2229929 (The NSF AI Institute for Artificial and Natural Intelligence). The views, opinions and/or findings expressed are those of the authors and should not be interpreted as representing the official views or policies of the National Science Foundation or the U.S. Government.
@@ Appendix A Psychology Theories
\__nicematrix_patch_booktabs:\__nicematrix_revert_colortbl: